# Fuzzy K-Means Clustering without Cluster Centroids

### v1

@@ Abstract.
Clustering, Fuzzy K-Means
1. Introduction
@@
Numerous clustering algorithms have been developed, including spectral clustering (Ng et al., 2001; Yang et al., 2022), K-Means (MacQueen et al., 1967; Lu et al., 2024; Nie et al., 2023), Kernel K-Means (Kim et al., 2005; Wang et al., 2022; Ren et al., 2021), and Fuzzy K-Means (Dunn, 1973; Xu et al., 2016; Nie et al., 2022b; Wang et al., 2023), among others.
@@
Spectral clustering (Ng et al., 2001), a graph-based clustering method (Shi and Malik, 2000), transforms clustering into an optimal graph partitioning problem. It begins by decomposing the graph’s Laplacian matrix to derive a low-dimensional sample representation, followed by employing K-Means to segment data points into K clusters. Similar to other graph-based clustering methods, spectral clustering involves constructing a similarity adjacency matrix among samples, performing eigendecomposition on this matrix, and then clustering the resulting low-dimensional representation. Although graph-based approaches often yield superior results, constructing the adjacency matrix and the instability of eigendecomposition are notable drawbacks.
@@
In contrast, K-Means clustering (MacQueen et al., 1967) is renowned for its simplicity and has been a staple method since its inception. This algorithm partitions a dataset into K subsets, iteratively updating cluster centroids and reassigning clusters to minimize the sum of distances between samples and their respective centroids, thereby achieving distinct and compact clustering. Extensions of K-Means, like K-Means++ (Arthur and Vassilvitskii, 2007), address the randomness in initial centroid selection. K-Means++ strategically selects initial centroids to enhance the likelihood of choosing representative sample points. For datasets not amenable to linear separation, Kernel K-Means (Kim et al., 2005; Wang et al., 2022) employs kernel methods to map data into a high-dimensional space, facilitating linear separability. The choice of kernel function in this context is critical and should reflect the data characteristics and problem complexity. Multi-kernel K-Means (Ren et al., 2021) integrates multiple kernel functions, proving effective in clustering complex, high-dimensional data. Additionally, K-Medoids (Ren et al., 2022; Newling and Fleuret, 2017) offers a robust alternative to traditional K-Means, using the most centrally located object in a cluster as the cluster center, thereby reducing the influence of outliers.
@@
The spectral clustering and K-Means clustering mentioned above, along with their related extension algorithms, are categorized as hard clustering. This approach strictly assigns each object to a specific cluster, following an ”either-or” principle. However, many real-world datasets exhibit fuzzy rather than clear category boundaries, a scenario where hard clustering algorithms struggle due to their inability to handle such fuzziness. To address this, the Fuzzy K-Means clustering algorithm (Dunn, 1973; Xu et al., 2016; Nie et al., 2022b; Wang et al., 2023) was introduced. It incorporates fuzzy theory to determine the degree of membership of each sample to all cluster centers by optimizing the objective function. This method’s strength lies in its capacity to express intermediate sample memberships, thereby capturing the uncertainty inherent in sample categorization and more accurately reflecting real-world scenarios. Enhanced versions of the original Fuzzy K-Means clustering have been developed, such as the Maximum Entropy Fuzzy Clustering algorithm (Zhou et al., 2016; Li et al., 2008), which constructs a bipartite graph based on information entropy. Furthermore, Par et al. introduced a hybrid Possibilistic Fuzzy K-Means algorithm (Zhou et al., 2016; Tang et al., 2019), combining Fuzzy K-Means with Possibilistic K-Means and applying a more relaxed constraint on the membership matrix.
@@
Despite these advancements, the algorithmic process of Fuzzy K-Means clustering mirrors that of traditional K-Means, requiring random initialization of cluster centroids and their update through weighted mean calculation. This dependency significantly impacts the performance and robustness of clustering results. Attempts to eliminate the influence of cluster centers in K-Means clustering have led to innovations like Ding’s (Ding and He, 2005) proof of the equivalence of non-negative matrix factorization to kernel K-Means clustering with relaxed orthogonal relationships. This method calculates clustering labels through matrix factorization. Additionally, connections between kernel K-Means and spectral clustering have been established (Dhillon et al., 2004). However, these methods necessitate post-processing, which can compromise clustering quality and deviate from the original problem’s solution. Centerless clustering, as proposed by Pei et al. (Pei et al., 2023a), minimizes the sum of intra-cluster distances without considering cluster centroids. Yet, its assumption of equal sample numbers in each cluster limits its effectiveness on real-world datasets with cluster imbalances. Currently, no research directly addresses clustering without considering cluster centroids in the context of Fuzzy K-Means.
@@ 2. Related Works
2.1. K-Means Clustering
K-Means clustering (MacQueen et al., 1967) partitions a given sample set. This method segregates the data set into K distinct subsets, with each subset representing a cluster. Given an input data matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$, where $N$ denotes the number of samples and $d$ represents the dimensionality of each sample, the goal of K-Means clustering is to minimize the following objective function:
(1) $\mathop{\min}\limits_{\textbf{X}_{1},\ldots,\textbf{X}_{K}}{\sum\limits_{l=1}^{K}\sum\limits_{\textbf{x}_{i}\in\textbf{X}_{l}}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$
@@ 2.1. K-Means Clustering
(2) $\mathop{\min}\limits_{\textbf{F}\in Ind,\textbf{U}}{\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}f_{il}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$
@@
2.2. Fuzzy K-Means Clustering
@@
2.2.1. Fuzzy K-Means Clustering
Fuzzy K-Means clustering (Dunn, 1973) extends the traditional K-Means clustering algorithm by introducing fuzzy parameters. In Fuzzy K-Means, each data point can belong to multiple cluster centroids, rather than being strictly assigned to a single cluster centroid as in traditional hard clustering. This inclusion of fuzzy parameters allows Fuzzy K-Means to more effectively handle uncertainties and fuzziness in data, thus offering a more accurate reflection of the inherent structure and relationships within the dataset.
@@ 2.2.1. Fuzzy K-Means Clustering
(3) $\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}^{r}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$
@@
(4) $y_{il}=\frac{\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{-\frac{2}{r-1}}}{\sum\nolimits_{j=1}^{K}\left(\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{-\frac{2}{r-1}}\right)}$
(5) $\textbf{u}_{l}=\frac{\sum\nolimits_{i=1}^{N}u_{il}^{r}\textbf{x}_{i}}{\sum\nolimits_{i=1}^{N}u_{il}^{r}}$
The Fuzzy K-Means clustering algorithm also requires the selection of K initial cluster centroids. It updates the membership degree matrix and cluster centroids iteratively. The stopping criteria for iterative FCM algorithms is given by $\max\left|y_{il}^{(t)}-y_{il}^{(t-1)}\right|<\epsilon$, where $\epsilon$ is a small positive number, and $t$ denotes the number of iterations (Bezdek, 1981).
2.2.2. Robust and Sparse Fuzzy K-Means Clustering
The performance of Fuzzy K-Means can be enhanced through regularization techniques (Namkoong et al., 2010). Xu et al. proposed an advanced variant, the Robust and Sparse Fuzzy K-Means clustering (RSFKM) (Xu et al., 2016), which integrates robustness and sparsity into the clustering process. The objective function of RSFKM is given by:
(6) $\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}\hat{d}_{il}+\lambda\|\textbf{Y}\|_{F}^{2}}$
@@ 2.2.2. Robust and Sparse Fuzzy K-Means Clustering
(7) $\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}+\lambda\|\textbf{Y}\|_{F}^{2}}$
@@
3. Methodology
3.1. Problem Formulation and objective
Fuzzy K-Means clustering is a widely utilized soft clustering algorithm. Unlike traditional K-Means clustering and other hard clustering methods, Fuzzy K-Means offers distinct advantages. It permits data points to be part of multiple clusters, each with varying degrees of membership. This flexibility makes it more adept at handling overlapping data, delivering more nuanced clustering outcomes. However, akin to K-Means, it confronts challenges, notably sensitivity to initial cluster centroid selection and vulnerability to outliers, owing to its reliance on centroid computation. To address these issues, we propose a modified approach: Fuzzy K-Means clustering without cluster centroids (FKMWC). FKMWC obviates the need for selecting and updating cluster centroids. The objective function of FKMWC is delineated below:
(8) $\mathop{\textrm{min}}\limits_{\textbf{Y}}\textrm{tr}(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y}\textbf{P}^{-1})+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$
@@ 3.1. Problem Formulation and objective
where $\mathbf{P}$ is a diagonal matrix and $\mathbf{D}$ is the distance matrix. If the matrix $\mathbf{D}$ in Eq. (8) represents the squared Euclidean distance, that is, $d_{ij}=\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|_{2}^{2}$, then Eq. (8) becomes equivalent to Eq. (7) as per Lemma 3.1. This implies that the objective of model (8) aligns with that of the Fuzzy $K$-Means Clustering.
@@
Lemma 0 ().
(9) $\mathop{\min}\limits_{\textbf{Y},\textbf{U}}\sum\limits_{i}\sum\limits_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}\Leftrightarrow\mathop{\min}\limits_{\textbf{Y}}\textrm{tr}(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y}\textbf{P}^{-1})$
Proof.
Let $J=\sum_{i}\sum_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}$, and then $\frac{\partial J}{\partial\textbf{u}_{j}}=2\sum\nolimits_{i}y_{ij}\textbf{u}_{j}-2\sum\nolimits_{i}\textbf{x}_{i}y_{ij}$. Then $\textbf{u}_{j}$ can be solved by taking $\frac{\partial J}{\partial\textbf{u}_{j}}=0$ because $J$ is a convex function. Thus we have the following
(10) $\textbf{u}_{j}=\frac{\sum\nolimits_{i}\textbf{x}_{i}y_{ij}}{\sum\nolimits_{i}y_{ij}}=\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\hat{p}_{j}^{-1}$
where $\hat{p}_{j}=\sum\nolimits_{i}y_{ij}$ and $\textbf{y}_{:j}$ is the $j$-th column of Y. Substituting Eq. (10) into $J$, and then it is clear that
(11) $J$ $=\sum\nolimits_{i}\sum\nolimits_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}$
$=\sum\nolimits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}\sum\nolimits_{j}y_{ij}+\sum\nolimits_{j}\textbf{u}_{j}\textbf{u}^{\textrm{T}}_{j}\sum\nolimits_{i}y_{ij}-2\sum\nolimits_{j}\textbf{u}_{j}\sum\nolimits_{i}\textbf{x}_{i}^{\textrm{T}}y_{ij}$
$=\sum\nolimits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}+\sum\nolimits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}-2\sum\nolimits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}$
$=\sum\nolimits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}-\sum\nolimits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}$
$=\textrm{tr}(\textbf{X}^{\textrm{T}}\textbf{X})-\textrm{tr}(\textbf{X}^{\textrm{T}}\textbf{Y}\textbf{P}^{-1}\textbf{Y}^{\textrm{T}}\textbf{X})$
$=\textrm{tr}\left(\textbf{X}^{\textrm{T}}(\textbf{I}-\textbf{Y}\textbf{P}^{-1}\textbf{Y}^{\textrm{T}})\textbf{X}\right)$
[…]
(12) $J$ $=\textrm{tr}\left(\textbf{X}^{\textrm{T}}(\textbf{I}-\textbf{S})\textbf{X}\right)$
$=\sum\nolimits_{i}\sum\nolimits_{j}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}$
$=\sum\nolimits_{i}\sum\nolimits_{j}d_{ij}<\textbf{f}_{i},\textbf{f}_{j}>$
$=\textrm{tr}(\textbf{F}^{\textrm{T}}\textbf{D}\textbf{F})$
$=\textrm{tr}(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y}\textbf{P}^{-1})$
Now Lemma 3.1 is proven. ∎

### v2

# Distance-based Fuzzy K-means Clustering without Cluster Centroids

@@ Abstract
Index Terms:
Clustering, Fuzzy K-Means, unsupervised learning.
1 Introduction
@@
Numerous clustering algorithms have been developed, including spectral clustering [1, 2], K-Means [3, 4, 5], Kernel K-Means [6, 7, 8], and Fuzzy K-Means [9, 10, 11, 12], among others.
@@
Spectral clustering [1], a graph-based clustering method [13], transforms clustering into an optimal graph partitioning problem. It begins by decomposing the graph’s Laplacian matrix to derive a low-dimensional sample representation, followed by employing K-Means to segment data points into K clusters. Similar to other graph-based clustering methods, spectral clustering involves constructing a similarity adjacency matrix among samples, performing eigendecomposition on this matrix, and then clustering the resulting low-dimensional representation. Although graph-based approaches often yield superior results, constructing the adjacency matrix and the instability of eigendecomposition are notable drawbacks.
@@
In contrast, K-Means clustering [3] is renowned for its simplicity and has been a staple method since its inception. This algorithm partitions a dataset into K subsets, iteratively updating cluster centroids and reassigning clusters to minimize the sum of distances between samples and their respective centroids, thereby achieving distinct and compact clustering. Extensions of K-Means, like K-Means++ [14], address the randomness in initial centroid selection. K-Means++ strategically selects initial centroids to enhance the likelihood of choosing representative sample points. For datasets not amenable to linear separation, Kernel K-Means [6, 7] employs kernel methods to map data into a high-dimensional space, facilitating linear separability. The choice of kernel function in this context is critical and should reflect the data characteristics and problem complexity. Multi-kernel K-Means [8] integrates multiple kernel functions, proving effective in clustering complex, high-dimensional data. Additionally, K-Medoids [15, 16] offers a robust alternative to traditional K-Means, using the most centrally located object in a cluster as the cluster center, thereby reducing the influence of outliers.
@@
The spectral clustering and K-Means clustering mentioned above, along with their related extension algorithms, are categorized as hard clustering. This approach strictly assigns each object to a specific cluster, following an ”either-or” principle. However, many real-world datasets exhibit fuzzy rather than clear category boundaries, a scenario where hard clustering algorithms struggle due to their inability to handle such fuzziness. To address this, the Fuzzy K-Means clustering algorithm [9, 10, 11, 12] was introduced. It incorporates fuzzy theory to determine the degree of membership of each sample to all cluster centers by optimizing the objective function. This method’s strength lies in its capacity to express intermediate sample memberships, thereby capturing the uncertainty inherent in sample categorization and more accurately reflecting real-world scenarios. Enhanced versions of the original Fuzzy K-Means clustering have been developed, such as the Maximum Entropy Fuzzy Clustering algorithm [17, 18], which constructs a bipartite graph based on information entropy. Furthermore, Par et al. introduced a hybrid Possibilistic Fuzzy K-Means algorithm [17, 19], combining Fuzzy K-Means with Possibilistic K-Means and applying a more relaxed constraint on the membership matrix.
@@
Despite these advancements, the algorithmic process of Fuzzy K-Means clustering mirrors that of traditional K-Means, requiring random initialization of cluster centroids and their update through weighted mean calculation. This dependency significantly impacts the performance and robustness of clustering results. Attempts to eliminate the influence of cluster centers in K-Means clustering have led to innovations like Ding’s [20] proof of the equivalence of non-negative matrix factorization to kernel K-Means clustering with relaxed orthogonal relationships. This method calculates clustering labels through matrix factorization. Additionally, connections between kernel K-Means and spectral clustering have been established [21]. However, these methods necessitate post-processing, which can compromise clustering quality and deviate from the original problem’s solution. Centerless clustering, as proposed by Pei et al. [22], minimizes the sum of intra-cluster distances without considering cluster centroids. Yet, its assumption of equal sample numbers in each cluster limits its effectiveness on real-world datasets with cluster imbalances. Currently, no research directly addresses clustering without considering cluster centroids in the context of Fuzzy K-Means.
@@ 2 Related Works
2.1 K-Means Clustering
K-Means clustering [3] partitions a given sample set. This method segregates the data set into K distinct subsets, with each subset representing a cluster. Given an input data matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$, where $N$ denotes the number of samples and $d$ represents the dimensionality of each sample, the goal of K-Means clustering is to minimize the following objective function:
$\mathop{\min}\limits_{\textbf{X}_{1},\ldots,\textbf{X}_{K}}{\sum\limits_{l=1}^{K}\sum\limits_{\textbf{x}_{i}\in\textbf{X}_{l}}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$ (1)
@@ 2.1 K-Means Clustering
$\mathop{\min}\limits_{\textbf{F}\in Ind,\textbf{U}}{\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}f_{il}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$ (2)
@@
2.2 Fuzzy K-Means Clustering
@@
2.2.1 Fuzzy K-Means Clustering
Fuzzy K-Means clustering [9] extends the traditional K-Means clustering algorithm by introducing fuzzy parameters. In Fuzzy K-Means, each data point can belong to multiple cluster centroids, rather than being strictly assigned to a single cluster centroid as in traditional hard clustering. This inclusion of fuzzy parameters allows Fuzzy K-Means to more effectively handle uncertainties and fuzziness in data, thus offering a more accurate reflection of the inherent structure and relationships within the dataset.
@@ 2.2.1 Fuzzy K-Means Clustering
$\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}^{r}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}}$ (3)
@@
$y_{il}=\frac{\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{-\frac{2}{r-1}}}{\sum\limits_{j=1}^{K}\left(\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{-\frac{2}{r-1}}\right)}$ (4)
$\textbf{u}_{l}=\frac{\sum\limits_{i=1}^{N}u_{il}^{r}\textbf{x}_{i}}{\sum\limits_{i=1}^{N}u_{il}^{r}}$ (5)
The Fuzzy K-Means clustering algorithm also requires the selection of K initial cluster centroids. It updates the membership degree matrix and cluster centroids iteratively. The stopping criteria for iterative FCM algorithms is given by $\max\left|y_{il}^{(t)}-y_{il}^{(t-1)}\right|<\epsilon$, where $\epsilon$ is a small positive number, and $t$ denotes the number of iterations [23].
2.2.2 Robust and Sparse Fuzzy K-Means Clustering
The performance of Fuzzy K-Means can be enhanced through regularization techniques [24]. Xu et al. proposed an advanced variant, the Robust and Sparse Fuzzy K-Means clustering (RSFKM) [10], which integrates robustness and sparsity into the clustering process. The objective function of RSFKM is given by:
$\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}\hat{d}_{il}+\lambda\|\textbf{Y}\|_{F}^{2}}$ (6)
@@ 2.2.2 Robust and Sparse Fuzzy K-Means Clustering
$\mathop{\min}\limits_{\textbf{Y},\textbf{U}}$ ${\sum\limits_{i=1}^{N}\sum\limits_{l=1}^{K}y_{il}\left\|\textbf{x}_{i}-\textbf{u}_{l}\right\|_{2}^{2}+\lambda\|\textbf{Y}\|_{F}^{2}}$ (7)
@@
3 Methodology
3.1 Problem Formulation and objective
Fuzzy K-Means clustering is a widely utilized soft clustering algorithm. Unlike traditional K-Means clustering and other hard clustering methods, Fuzzy K-Means offers distinct advantages. It permits data points to be part of multiple clusters, each with varying degrees of membership. This flexibility makes it more adept at handling overlapping data, delivering more nuanced clustering outcomes. However, akin to K-Means, it confronts challenges, notably sensitivity to initial cluster centroid selection and vulnerability to outliers, owing to its reliance on centroid computation. To address these issues, we introduce Lemma 1, which re-expresses Fuzzy K-Means from the perspective of manifold learning.
Lemma 1
$\mathop{\min}\limits_{\textbf{Y},\textbf{U}}\sum\limits_{i}\sum\limits_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}\Leftrightarrow\mathop{\min}\limits_{\textbf{Y}}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}$ (8)
where the manifold structure $\textbf{S}=\textbf{Y}\textbf{P}^{-1}\textbf{Y}^{\textrm{T}}$ represents the cluster structure in the data. P is a diagonal matrix and $p_{jj}=\hat{p}_{j}=\sum\limits_{i}y_{ij}$.
Proof 1
Let $J=\sum\limits_{i}\sum\limits_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}$, and then $\frac{\partial J}{\partial\textbf{u}_{j}}=2\sum\limits_{i}y_{ij}\textbf{u}_{j}-2\sum\limits_{i}\textbf{x}_{i}y_{ij}$. Then $\textbf{u}_{j}$ can be solved by taking $\frac{\partial J}{\partial\textbf{u}_{j}}=0$ because $J$ is a convex function. Thus we have the following
$\textbf{u}_{j}=\frac{\sum\limits_{i}\textbf{x}_{i}y_{ij}}{\sum\limits_{i}y_{ij}}=\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\hat{p}_{j}^{-1}$ (9)
where $\hat{p}_{j}=\sum\limits_{i}y_{ij}$ and $\textbf{y}_{:j}$ is the $j$-th column of Y. Substituting Eq. (9) into $J$, and then it is clear that
$\sum\limits_{i}\sum\limits_{j}y_{ij}\left\|\textbf{x}_{i}-\textbf{u}_{j}\right\|_{2}^{2}$ (10)
$\quad\quad=\sum\limits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}\sum\limits_{j}y_{ij}+\sum\limits_{j}\textbf{u}_{j}\textbf{u}^{\textrm{T}}_{j}\sum\limits_{i}y_{ij}-$
$\quad\quad\quad\quad 2\sum\limits_{j}\textbf{u}_{j}\sum\limits_{i}\textbf{x}_{i}^{\textrm{T}}y_{ij}$
$\quad\quad=\sum\limits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}+\sum\limits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}-$
$\quad\quad\quad\quad 2\sum\limits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}$
$\quad\quad=\sum\limits_{i}\textbf{x}_{i}\textbf{x}_{i}^{\textrm{T}}-\sum\limits_{j}\textbf{y}_{:j}^{\textrm{T}}\textbf{X}\textbf{X}^{\textrm{T}}\textbf{y}_{:j}\hat{p}_{j}^{-1}$
$\quad\quad=\textrm{tr}(\textbf{X}^{\textrm{T}}\textbf{X})-\textrm{tr}(\textbf{X}^{\textrm{T}}\textbf{Y}\textbf{P}^{-1}\textbf{Y}^{\textrm{T}}\textbf{X})$
$\quad\quad=\textrm{tr}\left(\textbf{X}^{\textrm{T}}(\textbf{I}-\textbf{Y}\textbf{P}^{-1}\textbf{Y}^{\textrm{T}})\textbf{X}\right)$
[…]
$J$ $=\textrm{tr}\left(\textbf{X}^{\textrm{T}}(\textbf{I}-\textbf{S})\textbf{X}\right)=\sum\limits_{i}\sum\limits_{j}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}$ (11)
Now Lemma 1 is proven.
Our approach constructs a manifold structure S from label Y, thereby ensuring label consistency for samples on the same manifold. This also eliminates the need to estimate the centroid matrix. By Lemma 1, Eq. (7) can be equivalent to Eq. (12). The objective function of our FKMWC is delineated below:
$\mathop{\textrm{min}}\limits_{\textbf{Y}}\sum\limits_{i}\sum\limits_{j}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$ (12)
$s.t.\quad s_{jj}=\sum_{i}{y_{ij}},\textbf{Y}\textbf{1}=\textbf{1},\textbf{Y}\geq\textbf{0}$
In order to facilitate the solution, we convert Eq. (12) into the form of Eq. (13) for solving
$\sum\limits_{i}\sum\limits_{j}s_{ij}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2}+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$ (13)
$\quad\quad=\sum\limits_{i}\sum\limits_{j}d_{ij}<\textbf{f}_{i},\textbf{f}_{j}>+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$
$\quad\quad=\textrm{tr}(\textbf{F}^{\textrm{T}}\textbf{D}\textbf{F})+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$
$\quad\quad=\textrm{tr}(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y}\textbf{P}^{-1})+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$
@@ 3.1 Problem Formulation and objective
where $\mathbf{P}$ is a diagonal matrix and $\mathbf{D}$ is the distance matrix. If the matrix $\mathbf{D}$ in Eq. (13) represents the squared Euclidean distance, that is, $d_{ij}=\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|_{2}^{2}$, then Eq. (13) becomes equivalent to Eq. (7) as per Lemma 1. This implies that the objective of model (13) aligns with that of the Fuzzy $K$-Means Clustering.